# Designing Parallel Algorithms for Community Detection using Arachne

### v1

@@ Abstract
The rise of graph data in various fields calls for efficient and scalable community detection algorithms. In this paper, we present parallel implementations of two widely used algorithms: Label Propagation and Louvain, specifically designed to leverage the capabilities of Arachne which is a Python-accessible, open-source framework for large-scale graph analysis. Our implementations achieve substantial speedups over existing Python-based tools like NetworkX and igraph, which lack efficient parallelization, and are competitive with parallel frameworks such as NetworKit. Experimental results show that Arachne-based methods outperform these baselines, achieving speedups of up to 710x over NetworkX, 75x over igraph, and 12x over NetworKit. Additionally, we analyze the scalability of our implementation under varying thread counts, demonstrating how different phases contribute to overall performance gains of the parallel Louvain algorithm. Arachne, including our community detection implementation, is open-source and available at https://github.com/Bears-R-Us/arkouda-njit.
@@
Graph Algorithms, Community Detection, Parallel Algorithms, High-Performance Data Analytics
@@ I Introduction
Community detection, also referred to as graph clustering, aims to partition the vertices of a network into groups characterized by dense intra-cluster connections and sparse inter-cluster connections. It is important to comprehend the structure of complex networks, as community members usually exhibit similar roles within the network. In recent years, community detection has had a wide range of applications in various fields, including urban transportation [1], social network [2], biology [3], FinTech [4], and more.
As networks continue to grow in size, it is necessary to utilize effective parallel algorithms. For graphs with millions to billions of edges, only (near) linear time community detection algorithms are practical. To address this, we focus on two parallel algorithms. Label Propagation [5] is a simple yet effective method where vertices iteratively adopt the most frequent community label among their neighbors until stable communities emerge. The Louvain algorithm [6], widely recognized for its heuristic modularity optimization, has proven highly efficient for clustering large, real-world networks. Its performance has been further enhanced through various optimizations and parallel implementations [7, 8, 9, 10, 11]. In this paper, we implement in parallel and evaluate these two algorithms. Furthermore, they are integrated into the open-source framework Arachne [12], enabling scalable exploratory graph analysis.
Arachne [12] is an actively developed framework, integrated with Arkouda [13], which combines the productivity of Python at the front end with the high-performance computing capability of the Chapel programming language [14] at the back end. Arachne provides a comprehensive set of graph analytics algorithms, organized within a cohesive framework, with all Chapel and Python code available on GitHub. Our research focuses on developing and integrating parallel graph algorithms into Arachne, enabling data scientists to take advantage of Python and use their laptops to conduct real-world graph analysis on large compute platforms productively.
@@
1) Scalable Parallel Implementations of Community Detection Algorithms: We develop high-performance parallel versions of two widely used community detection algorithms: Label Propagation and Louvain, optimized for large-scale graphs. By leveraging the parallel computing capabilities of Chapel, our implementations efficiently process graphs with billions of edges, achieving up to 710× speedup over NetworkX, 75× over igraph, and 12× over NetworKit.
2) Comprehensive Evaluation across Various Graphs: We conduct extensive experiments to assess the efficiency, scalability, and output quality of our implementations. Results show that our methods consistently outperform widely used Python-based libraries across a variety of real-world graphs, while delivering comparable or better solution quality than both sequential and parallel baselines.
3) Integration with the Arachne Graph Analytics Framework: This work demonstrates how Arachne, a flexible and scalable open-source framework, facilitates the efficient implementation and execution of advanced community detection algorithms. By leveraging Arachne’s design and capabilities, we enable seamless scalability and high performance, providing a practical tool for analyzing massive graphs. This integration highlights Arachne’s potential as a robust platform for a broad range of graph analytics tasks beyond community detection.
@@ II-A Notation
Let $G(V,E,w)$ be an undirected, weighted graph, where $V$ represents the set of $n$ vertices, $E\subseteq\{\{u,v\}\mid u,v\in V\}$ denotes the set of $m$ edges. $N(v)$ denotes the neighbor of $v$. An edge $e$ is said to be incident to a vertex $v$ if $v$ is one of the two vertices that $e$ connects. Loops, edges that connect a vertex to itself, are also allowed. The weight associated with an edge $e_{ij}$ is denoted by $w_{ij}$. In the case of an unweighted graph, we assume all edges have unit weight, i.e., $w_{ij}=1$. Following the convention of [15], the degree of a vertex is defined as the number of its neighbors:
@@
For the weighted degree, we sum the weights of all incident edges
@@
where $L_{v}$ is the weight of the loop on $v$ if one exists, and is 0 otherwise. Loops are counted twice, similar to how they are treated in directed graphs. The weighted volume of a set of vertices $S$ is the sum of their weighted degrees:
$vol_{w}(S)=\sum_{v\in S}\deg_{w}(v)$
@@
$cut_{w}(A,B)=\sum_{\{u,v\}\mid\{u,v\}\in E,u\in A,v\in B}w(\{u,v\})$
We use $cut_{w}(A)$ as shorthand for the cut between $A$ and the rest of the graph, i.e., $cut_{w}(A,V\setminus A)$. Similarly, $cut_{w}(v,A)$ is used to represent $cut_{w}(\{v\},A)$.
@@ II-B Community Detection
Community detection is a widely used technique in network science to partition a network into a set of disjoint communities, $C={c_{1},\dots,c_{n}}$, where each community $c_{i}\subseteq V$, $\cup c_{i}=V$, and $c_{i}\cap c_{j}=\emptyset$. While some methods allow for overlapping communities, in this paper, we assume communities are disjoint, so overlapping is not considered in this context.
@@ II-C Modularity
Modularity, introduced by Newman and Girvan in 2004 [16], is a widely used metric for evaluating the quality of communities identified by heuristic-based community detection algorithms. Communities with high modularity are characterized by dense internal connections and sparse connections between them. The modularity ranges from -0.5 to 1 [17], with higher values indicating better community structures. As mentioned by [18], modularity can be written as:
$\sum_{c\in C}\frac{vol_{w}(c)-cut_{w}(C)}{vol_{w}(V)}-\frac{vol_{W}(C)^{2}}{vol_{w}(V)^{2}}$
Modularity can be changed by moving vertices between communities. A move consists of removing a vertex from its current community and assigning it to a different one. Let $A$ represent $v$’s current community and $B$ the target community. Denote by $A^{-}$ and $B^{-}$ the communities $A$ and $B$ excluding $v$, respectively. The modularity change, $\triangle{Q}{v\to B}$, resulting from such a move is given by:
[…]
@@ II-D Arachne
Arachne is an extension to the existing library, Arkouda, which provides Pandas- and NumPy-like operations at a supercomputing scale. Arachne integrates into typical exploratory data analytics workflows within Arkouda by providing methods to automatically convert tabular data into graph structures, enabling large-scale graph operations and queries. It efficiently executes graph analytical kernels on shared memory and distributed memory, transferring data as needed between compute nodes and processors as required.
@@ III-A1 Algorithm
The Label Propagation Algorithm (LPA), introduced by Raghavan et al. [5], is a widely used method for community detection in large-scale networks due to its simplicity and scalability. The algorithm initializes each vertex with a unique label and iteratively updates vertex labels based on their neighbors. In each iteration, a vertex adopts the label with the highest aggregate edge weight in its neighborhood (for weighted graphs), effectively forming communities through local greedy updates. The process continues until convergence, defined as the number of label updates falling below a user-defined threshold $\theta$. LPA has near-linear time complexity; each iteration takes $\mathcal{O}\left(m\right)$ time, and the total runtime is $\mathcal{O}\left(L\cdot m\right)$, where $L$ is the number of iterations. Although $L$ depends on network structure rather than size, LPA typically converges in only a few iterations in practice, despite the lack of formal guarantees.
@@ III-A2 Implementation
1:Input: Graph $G=(V,E)$, $maxIteration$, $threshold$
@@
4:     $C\leftarrow[0..|V|]$triangleright initialized as singleton
5:     $V_{active}\leftarrow V$triangleright active vertices in each iteration
6:     $\Delta N\leftarrow 0$triangleright number of vertices changed
@@
8:         $\Delta N\leftarrow\textsc{PLPMove}(G,C)$
@@
15:function PLPMove($G,C$)
@@
17:     Parallel for active vertex $v\in V$ do
18:           $c^{*}\leftarrow argmax_{c}\{\sum_{u\in N(v):C(u)=c}w(v,u)\}$
@@
20:               $V_{active}\leftarrow V_{active}\setminus{v}$
@@
25:           $V_{active}\leftarrow V_{active}\cup N(v)$
26:     return $\Delta N$
27:end function
Algorithm 1 shows the pseudocode of our parallel label propagation implementation. The algorithm is parallelized by dividing the vertex set across threads, each operating independently on a shared label array (line 14). This asynchronous approach accelerates convergence but may lead to race conditions—when evaluating the neighborhood of vertex $u$ (line 15), its neighbor $v$ might hold either the previous label $C_{i-1}(v)$ or the current label $C_{i}(v)$. However, these race conditions are acceptable and even beneficial. In the original description of the algorithm [5], the vertices are traversed in random order. We make this step optional in our parallel version, relying instead on the inherent randomization provided by thread execution.
To improve efficiency, we maintain a set of active vertices, $V_{active}$, which is initialized in line 5. A vertex becomes inactive if its label remains unchanged after an iteration, as implemented in lines 19 and 20, and is removed from the active set. It is reactivated only if one of its neighbors updates its label, as handled in line 25. This selective update strategy significantly reduces redundant computations and focuses computation on regions of the graph still undergoing change. The algorithm terminates when the number of updated vertices falls below a user-defined threshold, which is checked in line 9, or when the predefined maximum number of iterations is reached, as specified in line 7. These termination criteria ensure efficient execution while preserving community quality, particularly in the presence of high-degree vertices that may otherwise delay convergence.
@@ III-B Parallel Louvain Algorithm
The algorithm proceeds through two iterative phases: the local-moving phase and the aggregation phase. During the local-moving phase, each vertex $v$ evaluates whether moving to a neighboring community would increase modularity, and greedily joins the community that provides the highest gain. In the aggregation phase, all vertices within the same community are merged into a super-vertex, reducing the graph’s size.
@@
In our parallel implementation of the Louvain algorithm, we go beyond the straightforward approach of parallelizing the main for-loop. Our primary optimizations leverage the features of Arachne and Arkouda to accelerate both the local-moving and aggregation phases of the algorithm. The specific improvements and detailed workings of our approach are outlined below:
@@ III-B1 Local-moving Phase
Algorithm 2 Parallel Local-Moving
1:Input: Graph $G=(V,E)$, $maxIteration$
@@
4:     $comID[v]\leftarrow vertexID[v]$triangleright communityID
5:     $volVertex[v]\leftarrow w(v)$triangleright volume of vertex
6:     $volCom[v]\leftarrow w(v)$triangleright volume of community
@@
25:      $needCheck\leftarrow tmpNeedCheck$
26:until No vertex changes community or iteration count
27:       exceeds $maxIteration$
28:return $comID$
@@
At the beginning of each local-moving phase, since we calculate $\triangle Q$ using Equation. 1, we need to initialize two distributed arrays in parallel as shown in 1 of Algorithm. 2: $volVertex$ and $volCom$, representing the volume of vertices and the volume of community. In line 4, each vertex is assigned its own community by setting the community ID to its vertex ID, effectively creating singleton communities. As a result, both $volVertex$ and $volCom$ are arrays initialized in lines 5-6 with lengths equal to the number of vertices, and their initial values are set to $w(v)$, the weight of the respective vertex.
The core of our local moving phase lies in the parallel vertex movement evaluation process starting from line 9. For each vertex $v$, the algorithm concurrently examines if moving $v$ to one of its neighboring communities yields a positive $\triangle Q$. Specifically in lines 13-15, we parallelly compute the $\triangle Q$ for each neighboring community. After identifying a neighboring community that yields the maximum increase in $\triangle Q$ in line 16, the algorithm updates both the community membership and volume information. When a beneficial move is found, lines 17-19 show how we adjust the volumes by simultaneously decreasing $volCom[c]$ and increasing $volCom[d]$ by $volVertex[v]$, followed by updating the vertex’s community ID.
To optimize computation, line 7 introduces an atomic boolean array, $needCheck$, which tracks whether a vertex should be reconsidered for movement in the next iteration. The parallel processing in line 11 ensures that a vertex is only evaluated if it’s marked for checking, which happens when either its community or that of one of its neighbors has changed. This selective processing significantly reduces unnecessary computations. The operation shown in line 25 updates the $needCheck$ array for the next iteration based on the changes made in the current round. As specified in line 26, the local-moving phase continues this parallel process until either no vertex changes its community or the maximum iteration count is reached.

### v2

@@ Abstract
The rise of graph data in various fields calls for efficient and scalable community detection algorithms. In this paper, we present parallel implementations of two widely used algorithms: Label Propagation and Louvain, specifically designed to leverage the capabilities of Arachne, which is a Python-accessible open-source framework for large-scale graph analysis. Our implementations achieve substantial speedups over existing Python-based tools like NetworkX and igraph, which lack efficient parallelization, and are competitive with parallel frameworks such as NetworKit. Experimental results show that Arachne-based methods outperform these baselines, achieving speedups of up to 710x over NetworkX, 75x over igraph, and 12x over NetworKit. Additionally, we analyze the scalability of our implementation under varying thread counts, demonstrating how different phases contribute to overall performance gains of the parallel Louvain algorithm. Arachne, including our community detection implementation, is open-source and available at https://github.com/Bears-R-Us/arkouda-njit.
@@
Graph Algorithms, High-Performance Computing, Data Science
@@ I Introduction
Community detection, also referred to as graph clustering, aims to partition the vertices of a network into groups characterized by dense intra-cluster connections and sparse inter-cluster connections. It is important to comprehend the structure of complex networks, as community members usually exhibit similar roles within the network. In recent years, community detection has had a wide range of applications in various fields, including urban transportation [1], social networks [2], biology [3], FinTech [4], and more.
As networks continue to grow in size, it is necessary to utilize effective parallel algorithms. For graphs with millions to billions of edges, only (near) linear-time community-detection algorithms are practical. To address this, we focus on two parallel algorithms. Label propagation [5] is a simple yet effective method in which vertices iteratively adopt the most frequent community label among their neighbors until stable communities emerge. The Louvain algorithm [6], widely recognized for its heuristic modularity optimization, has proven highly efficient for clustering large, real-world networks. Its performance has been further improved through various optimizations and parallel implementations [7, 8, 9, 10, 11]. In this paper, we implement these two algorithms in parallel and evaluate them. Furthermore, they are integrated into the Arachne open-source framework [12], allowing scalable exploratory graph analysis.
Arachne [12] is an actively developed framework, integrated with Arkouda [13], which combines the productivity of Python at the front end with the high-performance computing capability of the Chapel programming language [14] at the back end. Arachne provides a comprehensive set of graph analytics algorithms, organized within a cohesive framework, with all Chapel and Python code available on GitHub. Our research focuses on developing and integrating parallel graph algorithms into Arachne, enabling data scientists to take advantage of Python and use their laptops to perform real-world graph analysis on large computing platforms productively.
@@
1) Scalable Parallel Implementations of Community Detection Algorithms: We develop high-performance parallel versions of two widely used community detection algorithms: Label propagation and Louvain, optimized for large-scale graphs. Using the parallel computing capabilities of Chapel, our implementations efficiently process graphs with billions of edges, achieving up to 710× speedup over NetworkX, 75× over igraph, and 12× over NetworKit.
2) Comprehensive Evaluation across Various Graphs: We conduct extensive experiments to assess the efficiency, scalability, and output quality of our implementations. The results show that our methods consistently outperform widely used Python-based libraries across a variety of real-world graphs, while delivering comparable or better solution quality than both sequential and parallel baselines.
3) Integration with the Arachne Graph Analytics Framework: This work demonstrates how Arachne, a flexible and scalable open-source framework, facilitates the efficient implementation and execution of advanced community detection algorithms. By leveraging Arachne’s design and capabilities, we enable seamless scalability and high performance, providing a practical tool for analyzing massive graphs. This integration highlights Arachne’s potential as a robust platform for a broad range of graph-analysis tasks beyond community detection.
@@ II-A Notation
Let $G(V,E,w)$ be an undirected, weighted graph, where $V$ represents the set of $n$ vertices, $E\subseteq\{\{u,v\}\mid u,v\in V\}$ denotes the set of $m$ edges. $N(v)$ denotes the neighbor of $v$. An edge $e$ is said to be incident to a vertex $v$ if $v$ is one of the two vertices that connects $e$. Loops, edges that connect a vertex to itself, are also allowed. The weight associated with an edge $e_{ij}$ is denoted by $w_{ij}$. In the case of an unweighted graph, we assume that all edges have unit weight, i.e., $w_{ij}=1$. Following the convention of [15], the degree of a vertex is defined as the number of its neighbors:
@@
For the weighted degree, we sum the weights of all incident edges, defined as:
@@
where $L_{v}$ is the weight of the loop on $v$ if one exists and is 0 otherwise. Loops are counted twice, similar to how they are treated in directed graphs. The weighted volume of a set of vertices $S$ is the sum of their weighted degrees:
$\textit{vol}_{w}(S)=\sum_{v\in S}\deg_{w}(v)$
@@
$\textit{cut}_{w}(A,B)=\sum_{\{u,v\}\mid\{u,v\}\in E,u\in A,v\in B}w(\{u,v\})$
We use $cut_{w}(A)$ as shorthand for the cut between $A$ and the rest of the graph, that is, $cut_{w}(A,V\setminus A)$. Similarly, $cut_{w}(v,A)$ is used to represent $cut_{w}(\{v\},A)$.
@@ II-B Community Detection
Community detection is a widely used technique in network science to partition a network into a set of disjoint communities $C={c_{1},\dots,c_{n}}$, where each community $c_{i}\subseteq V$, $\cup c_{i}=V$, and $c_{i}\cap c_{j}=\emptyset$. Although some methods allow for overlapping communities, in this paper we assume communities to be disjoint, so overlapping is not considered in this context.
@@ II-C Modularity
Modularity, introduced by Newman and Girvan in 2004 [16], is a widely used metric to evaluate the quality of communities identified by heuristic-based community detection algorithms. Communities with high modularity are characterized by dense internal connections and sparse connections between them. The modularity ranges from -0.5 to 1 [17], with higher values indicating better community structures. As mentioned in [18], modularity can be written as:
$\sum_{c\in C}\frac{\textit{vol}_{w}(c)-\textit{cut}_{w}(C)}{\textit{vol}_{w}(V)}-\frac{vol_{w}(C)^{2}}{vol_{w}(V)^{2}}$
Modularity can be changed by moving vertices between communities. A move consists of removing a vertex from its current community and assigning it to a different one. Let $A$ represent the current community of $v$ and $B$ the target community. Denote by $A^{-}$ and $B^{-}$ the communities $A$ and $B$ that exclude $v$, respectively. The modularity change, $\triangle{Q}{v\to B}$, resulting from such a move is given by:
$\text{$\triangle{Q}_{v\to B}=2*(\frac{\textit{cut}_{w}(v,B^{-})-\textit{cut}_{w}(v,A^{-})}{\textit{vol}_{w}(V)}-$}\\ \text{$\deg_{w}(V)*\frac{\textit{vol}_{w}(B^{-})-\textit{vol}_{w}(A^{-})}{\textit{vol}_{w}(V)^{2}})$}$ (1)
@@ II-D Arachne
Arachne is an extension to the existing library, Arkouda, which provides Pandas- and NumPy-like operations at a supercomputing scale. Arachne integrates into typical exploratory data analytics workflows within Arkouda by providing methods to automatically convert tabular data into graph structures, enabling large-scale graph operations and queries. It efficiently executes graph analytical kernels in shared and distributed memory, transferring data as needed between compute nodes and processors as required.
@@ III-A1 Algorithm
The Label Propagation Algorithm (LPA), introduced by Raghavan et al. [5], is a widely used method for community detection in large-scale networks due to its simplicity and scalability. The algorithm initializes each vertex with a unique label and iteratively updates vertex labels based on their neighbors. In each iteration, a vertex adopts the label with the highest aggregate edge weight in its neighborhood (for weighted graphs), effectively forming communities through local greedy updates. The process continues until convergence, defined as the number of label updates falling below a user-defined threshold $\theta$. LPA has a near-linear time complexity; Each iteration takes $\mathcal{O}\left(m\right)$ time, and the total running time is $\mathcal{O}\left(L\cdot m\right)$, where $L$ is the number of iterations. Although $L$ depends on the network structure rather than the size, LPA typically converges in only a few iterations in practice, despite the lack of formal guarantees.
@@ III-A2 Implementation
1:Input: Graph $G=(V,E)$, maxIterationmaxIteration, thresholdthreshold
@@
4:  $C\leftarrow[0..|V|]$ $\triangleright$ initialized as singleton
5:  $V_{\textit{active}}\leftarrow V$ $\triangleright$ active vertices in each iteration
6:  $\Delta N\leftarrow 0$ $\triangleright$ number of vertices changed
@@
8:   $\Delta N\leftarrow\textsc{PLPMove}(G,C,V_{\textit{active}})$
@@
15:function PLPMove($G,C,V_{\textit{active}}$)
@@
17:  parallel for $v\in V_{\textit{active}}$ do
18:      $c^{*}\leftarrow\textit{argmax}_{c}\{\sum_{u\in N(v):C(u)=c}w(v,u)\}$
@@
20:       $V_{\textit{active}}\leftarrow V_{\textit{active}}\setminus{v}$
@@
25:      $V_{\textit{active}}\leftarrow V_{\textit{active}}\cup N(v)$
26:  end parallel for
27:  return $\Delta N$
28:end function
Algorithm 1 shows the pseudocode of our parallel label propagation implementation. The algorithm is parallelized by dividing the vertex set across threads, each operating independently on a shared label array (line 14). This asynchronous approach accelerates convergence, but can lead to race conditions: When evaluating the neighborhood of vertex $u$ (line 15), its neighbor $v$ could hold the previous label $C_{i-1}(v)$ or the current label $C_{i}(v)$. However, these race conditions are acceptable and even beneficial. In the original description of the algorithm [5], the vertices are traversed in random order. We make this step optional in our parallel version, relying instead on the inherent randomization provided by thread execution.
To improve efficiency, we maintain a set of active vertices, $V_{active}$, which is initialized at line 5. A vertex becomes inactive if its label remains unchanged after an iteration, as implemented in lines 19 and 20, and is removed from the active set. It is reactivated only if one of its neighbors updates its label, as handled in line 25. This selective update strategy significantly reduces redundant computations and focuses computation on regions of the graph that are still undergoing change. The algorithm terminates when the number of updated vertices falls below a user-defined threshold, which is checked in line 9, or when the predefined maximum number of iterations is reached, as specified in line 7. These termination criteria ensure efficient execution while preserving community quality, particularly in the presence of high-degree vertices that may otherwise delay convergence.
@@ III-B Parallel Louvain Algorithm
The algorithm proceeds through two iterative phases: the local-moving phase and the aggregation phase. During the local-moving phase, each vertex $v$ evaluates whether moving to a neighboring community would increase modularity and greedily joins the community that provides the highest gain. In the aggregation phase, all vertices within the same community are merged into a super-vertex, reducing the graph’s size.
@@
In our parallel implementation of the Louvain algorithm, we go beyond the straightforward approach of parallelizing the main for-loop. Our primary optimizations take advantage of the features of Arachne and Arkouda to accelerate both the local-moving and aggregation phases of the algorithm. The specific improvements and detailed workings of our approach are described below.
@@ III-B1 Local-moving Phase
Algorithm 2 Parallel Local-Moving Phase
1:Input: Graph $G=(V,E)$, maxIterationmaxIteration
@@
4:  $comID[v]\leftarrow vertexID[v]$ $\triangleright$ communityID
5:  $volVertex[v]\leftarrow w(v)$ $\triangleright$ volume of vertex
6:  $volCom[v]\leftarrow w(v)$ $\triangleright$ volume of community
@@
25:      end if
26:     end if
27:    end parallel for
28:    $needCheck\leftarrow tmpNeedCheck$
29:until No vertex changes community or iteration count
30:   exceeds maxIterationmaxIteration
31:return comIDcomID
@@
At the beginning of each local-moving phase, since we calculate $\triangle Q$ using Eq. 1, we need to initialize two arrays: volVertex and volCom, representing the volume of vertices and the volume of the community. In line 4, each vertex is assigned its own community by setting the community ID to its vertex ID, effectively creating singleton communities. As a result, both volVertex and volCom are arrays initialized in lines 5-6 with lengths equal to the number of vertices, and their initial values are set to $w(v)$, the weight of the respective vertex.
The core of our local moving phase lies in the parallel vertex movement evaluation process starting from line 9. For each vertex $v$, the algorithm concurrently examines whether moving $v$ to one of its neighboring communities yields a positive $\triangle Q$. Specifically in lines 13-15, we compute the $\triangle Q$ in parallel for each neighboring community. After identifying a neighboring community that produces the maximum increase in $\triangle Q$ on line 16, the algorithm updates both the community membership and volume information. When a beneficial move is found, lines 17-19 show how we adjust the volumes by simultaneously decreasing $\textit{volCom}[c]$ and increasing $\textit{volCom}[d]$ by $\textit{volVertex}[v]$, followed by updating the vertex’s community ID.
To optimize computation, line 7 introduces an atomic Boolean array, needCheck, which tracks whether a vertex should be reconsidered for movement in the next iteration. The parallel processing in line 11 ensures that a vertex is only evaluated if it is marked for checking, which happens when either its community or that of one of its neighbors has changed. This selective processing significantly reduces unnecessary computations. The operation shown on line 25 updates the $needCheck$ array for the next iteration based on the changes made in the current round. As specified in line 26, the local-moving phase continues this parallel process until either no vertex changes its community or the maximum iteration count is reached.
We note that, similar to other parallel Louvain implementations, the final result is not strictly deterministic. Due to concurrent updates or variations in thread scheduling, the final community assignment and the achieved modularity may differ across runs. This nondeterminism is a common characteristic of parallel heuristics and does not affect the overall quality or scalability of the algorithm [9].